# Simulating fermions with exponentially lower overhead

### v1

@@ IV Transformations between Encodings
2:   Perform an inorder traversal over the qubit nodes, $r\leftarrow\left\lceil N/2\right\rceilth$ node
3:   $q_{1},\dots,q_{h}\leftarrow$ path from root to $r$
4:   Apply a $\operatorname{CNOT}$ ladder to rotate $r$ up to become the root
5:   Apply balance-binary on the the bniary subtree rooted at $r$’s left child (if it exists)
6:   Apply balance-binary on the the bniary subtree rooted at $r$’s right child (if it exists)
7: Apply balance-binary for starting tree shape
8: Apply inverse circuit of balance-binary for target tree shape
Figure 15: In this example execution of Algorithm 1 on $N=11$ qubits, node $6$ is the middle node which should become the root. Then $\operatorname{CNOT}_{46}\operatorname{CNOT}_{69}\operatorname{CNOT}_{56}\operatorname{CNOT}_{67}$ rotates node 66 up to become the root. This $\operatorname{CNOT}$ circuit can be implemented in depth $O(\log N)$ via Lemma 4. The Majorana leaves are not drawn.
@@
There exists a circuit of depth $O(\log^{2}N)$, consisting of $\operatorname{CNOT}$ and $S$ gates, transforming between any two product-preserving ternary tree maps.
@@
To begin the implementation of Algorithm 2, as illustrated in Fig. 16, we first apply Algorithm 1 to transform the binary subtree of the root, as well as the binary subtree of every middle child node, into one with no left children. All of these subtrees are independent of each other, so their transformations can be all done in parallel in depth $O(\log^{2}N)$.
@@
We now have a ternary tree with the Jordan-Wigner shape that maps the vacuum state to $\left|0^{n}\right\rangle$. By [33, Lemma 5.9], all ternary tree mappings with the Jordan-Wigner shape mapping the vacuum state to $\left|0^{n}\right\rangle$ must be equivalent to the exact Jordan-Wigner mapping up to a fermionic permutation, pair braiding (paired Majoranas $(\gamma_{2k-1},\gamma_{2k})$ transforming to $(\gamma_{2k},-\gamma_{2k-1})$), and Pauli sign changes. The latter two may be fixed by constant depth single-qubit gates, leaving the fermion permutation as the only degree of freedom.
This means that, given any two product-preserving ternary-tree maps $\phi_{1}$ and $\phi_{2}$, there exist circuits ${C}_{1}$ and ${C}_{2}$ of depth $O(\log^{2}N)$ such that ${C}_{1}\phi_{1}=\phi_{\mathrm{JW}}U_{\sigma}$ and ${C}_{2}\phi_{2}=\phi_{\mathrm{JW}}U_{\tau}$ for some permutations $\sigma,\tau\colon[N]\to[N]$. (Recall from Eq. (5) that $\phi_{\mathrm{JW}}U_{f}=\Phi_{\mathrm{JW}}(U_{f})\phi_{\mathrm{JW}}$.) By Theorem 2, there exists a circuit D{D} that implements $\Phi_{\mathrm{JW}}(U_{\tau\sigma^{-1}})$ in depth $O(\log^{2}N)$. Then $C_{2}^{-1}DC_{1}$ is a circuit of depth $O(\log^{2}N)$ transforming from $\phi_{1}$ to $\phi_{2}$. ∎
@@ V Applications
In this section, we discuss applications of our methods to time evolution under widely studied fermionic models. Our results are summarized in Table 1, where we compare to the best previous methods that we are aware of.
We consider simulation of fermionic Hamiltonians of the form
@@
where $n_{i}=a^{\dagger}_{i}a_{i}$, and $J_{ij},U_{ij}$ are the hopping and interaction coefficient matrices, respectively.
@@
The fermionic fast Fourier transform diagonalizes the hopping Hamiltonian provided $J_{ij}$ is translationally invariant on a $d$-dimensional hypercubic lattice with periodic boundary conditions [2] (we focus on the 1-dimensional case here since the $d$-dimensional FFFT is simply the one-dimensional FFFT applied in all dimensions, just as for the usual FFT). Concretely,
@@
In the following, we show that the FFFT can be applied in polylog depth using our scheme, implying that time evolution under the hopping Hamiltonian for translationally-invariant systems can be done in polylog depth.
@@
Corollary 3.
Given any permutation $\sigma\colon[N]\to[N]$, there exists a circuit of depth $O(\log N)$ with $O(N)$ ancillas and measurement and feedback implementing $\Phi_{\mathrm{JW}}(U_{\sigma})$. In addition, there is a circuit of depth $O(\log N)$ implementing the FFFT using $O(N)$ ancillas and measurement and feedback.
@@
The materials Hamiltonian—or equivalently, quantum chemistry in the plane-wave dual basis [2]—is essentially a more complicated version of the Fermi-Hubbard model: the coefficients $J_{ij}$ and $U_{ij}$ are full rank, but both are translationally invariant [2]. Therefore, the hopping part can be compressed again to $O(\log^{2}N)$ depth without ancillas and $O(\log N)$ depth with ancillas by using the FFFT and our permutations. By contrast, because of the unboundedness of the hoppings, no scheme is known how to achieve this depth in the compact encoding. Instead, the depth in the compact encoding defaults to the same asymptotic depth as using fSWAP networks with transpositions [23] as the same geometric restrictions apply. A scheme to reduce the depth to $O(\log N)$ was proposed in Ref. [10], but required $O(N^{2})$ ancillas. For the interaction term, no scheme is known to apply it in polylog depth without ancillas; however, an $O(\log N)$-depth scheme is known using $O(N\log N)=\tilde{O}(N)$ ancillas [24]. Therefore, both of our schemes achieve polylog depth for a single Trotter step. Finally, for obtaining the end-to-end depth for simulating time evolution until time $T$, we need to multiply this by the number of Trotter steps. The best known upper bound is $O(NT)$ from Ref. [40], assumes arbitrarily large Trotter order and fixed total simulation error. Therefore, our schemes achieve $\tilde{O}(NT)$ depth complexity with $\tilde{O}(N)$ qubits. This quasilinear-in-number-of-plane-waves scaling of qubits and depth has to our knowledge not been achieved for qubit quantum computers yet and this scaling is on par (asymptotically) with fermionic quantum computers [10].
(Sub-)routine Compact FSWAP pFSWAP This work This work with anc.
@@
Fermion permut. NN NN NN 0 11 $N^{2}$ $\log^{2}N$ 0 $\log N$ NN
FFFT $\log N NN$ $N\log N$ 0 $\log N$ $N^{2}$ $\log^{2}N$ 0 $\log N$ NN
Materials (1 step) NN NN NN 0 $\log N$ $N^{2}$ $\log^{2}N$ $N\log N$ $\log N$ $N\log N$
Materials (full) $^{2} NN$ $N^{2}$ 0 $N\log N$ $N^{2}$ $N\log^{2}N$ $N\log N$ $N\log N$ $N\log N$
2D Fermi-Hubbard 11 NN $\sqrt{N}$ 0 $\log N$ $N^{2}$ $\log^{2}N$ 0 $\log N$ NN
Table 1: Asymptotic circuit depth and ancilla count for key fermionic subroutines and applications. All quantities are $O(\cdot)$ upper bounds. Compact refers to any compact encoding, e.g., Verstraete-Cirac [27] or Derby-Klassen [21]; FSWAP is the transposition-based fSWAP network from Ref. [23]; pFSWAP refers to parallel FSWAP from Ref. [10] using $O(N^{2})$ ancillas to brute-force parallelize the CZ circuits in the fSWAP networks; “2D Fermi-Hubbard” applies, more generally, to any model with $O(1)$-range hopping and interactions and assumes periodic boundary conditions for pFSWAP and our work (as we assume the usage of the FFFT for the kinetic term), and the gate counts are for a single Trotter step. The number of Trotter steps $O(N)$ used in “Materials (full)” assumes arbitrarily large Trotter order, fixed evolution time, and fixed simulation error. In the “this work with anc.” column, we also make use of ancillas in the permutation step, whereas in “this work” column, we only use ancillas for parallelizing the interaction part of the Hamiltonian [24].
@@ VI Discussion and Outlook
In this work, we exponentially reduce the worst-case circuit-depth overhead of time-evolving under arbitrary fermionic (parity-preserving) Hamiltonians on qubit-based quantum computers by introducing a new routing protocol that implements any fermionic permutation in the Jordan-Wigner encoding with depth $O(\log^{2}N)$ without ancillas and $O(\log N)$ when $O(N)$ ancillas and mid-circuit measurements are allowed. We also extend these results to arbitrary product-preserving ternary-tree encodings by introducing efficient conversions from those encodings and the Jordan-Wigner encoding. This advance removes the long-standing worst-case linear-depth overhead of fermion-to-qubit mappings, enabling key primitives such as the fermionic fast Fourier transform (FFFT) and single-Trotter-step material simulation to run in polylogarithmic depth using only $N$ qubits for $N$ fermionic modes.
Our analysis assumes hardware with all-to-all connectivity. However, our results also imply improvements for the limited-connectivity case: the worst-case overhead of fermionic simulation in limited-connectivity qubit systems is $O(\log^{2}N)$ times the worst-case depth to perform a qubit permutation. For example, a qubit system with grid connectivity could implement the FFFT and other time-evolution primitives in depth $O(\sqrt{N}\log^{2}N)$ using our protocol. Likewise, current neutral-atom systems could implement the FFFT and other time-evolution primitives in depth $O(\sqrt{N}\log^{3}N)$, or depth $O(\log^{3}N)$ with an additional spatial light modulator array [26]. We contrast this with the prior best known FFFT algorithm, which achieved $O(N\log N)$ depth [2] on a linear connectivity graph, but with no possibility to improve if better-than-linear connectivity is available.
We emphasize that while protocols for simulation of specific Hamiltonians with both low qubit connectivity and simulation overhead were previously known (see e.g. [27, 21]), this work tightens the best known general upper bound for time-evolving under any fermionic (parity-preserving) Hamiltonian.
Our work motivates future study in the reduction of fermionic simulation overhead in limited-connectivity qubit systems both for general and specific Hamiltonians. In addition, our improvements motivate further investigations into the tradeoff of ancilla count, mid-circuit measurement and feedback, and depth overhead.
It would also be illuminating to go beyond considering the asymptotic scaling of the depth overhead of the fermion-to-qubit mapping. In particular, for practical applications, it will be crucial to explicitly compare the gate counts and gate depths, including all prefactors, between our approach and previous ones.
Finally, it may be fruitful to consider whether similar circuit manipulations could be used to improve boson-to-qubit encodings [41, 42, 43].
Note added.
During the final stages of completing this manuscript, Maskara et al. posted a preprint [44] describing a scheme achieving $O(\log N)$ depth overhead in the Jordan-Wigner encoding using efficient permutations, employing $O(N)$ ancillas and feedforward of intermediate measurement results in order to compress CNOT ladders.

### v2

@@ IV Transformations between Encodings
2:   while some node has $\operatorname{hd}\geq 6$ do
3:    for all nodes AA with depth divisible by 33 do
4:      if $\operatorname{hd}(A)\geq 4$ then
5:       $A^{\prime},A^{\prime\prime},A^{\prime\prime\prime}\leftarrow$ first three nodes in heavy path from AA
6:       Apply $\operatorname{CNOT}$ rotations according to Table 1
7: Apply balance-binary to the starting tree
8: Apply Algorithm 1 on the resulting tree
@@
There exists a circuit of depth $O(\log^{2}N)$, consisting of $\operatorname{CNOT}$ and $S$ gates, transforming between any two product-preserving ternary tree mappings.
@@
To begin the implementation of Algorithm 3, as illustrated in Fig. 15, we first apply Algorithm 2 to transform the binary subtree of the root, as well as the binary subtree of every middle child node, into one with no left children. All of these subtrees are independent of each other, so all their transformations can be done in parallel in depth $O(\log N)$.
@@
We now have a ternary tree with the Jordan-Wigner shape that maps the vacuum state to $\left|0^{n}\right\rangle$. By [25, Lemma 5.9], all ternary tree mappings with the Jordan-Wigner shape mapping the vacuum state to $\left|0^{n}\right\rangle$ must be equivalent to the exact Jordan-Wigner mapping up to a fermionic permutation, pair braiding (paired Majoranas $(\gamma_{2k-1},\gamma_{2k})$ transforming to $(\gamma_{2k},-\gamma_{2k-1})$), and Pauli sign changes. The latter two may be fixed by constant-depth single-qubit gates, leaving the fermion permutation as the only degree of freedom.
This means that, given any two product-preserving ternary-tree mappings $\phi_{1}$ and $\phi_{2}$, there exist circuits ${C}_{1}$ and ${C}_{2}$ of depth $O(\log^{2}N)$ such that ${C}_{1}\phi_{1}=\phi_{\mathrm{JW}}U_{\sigma}$ and ${C}_{2}\phi_{2}=\phi_{\mathrm{JW}}U_{\tau}$ for some permutations $\sigma,\tau\colon[N]\to[N]$. (Recall from Eq. (3) that $\phi_{\mathrm{JW}}U_{f}=\Phi_{\mathrm{JW}}(U_{f})\phi_{\mathrm{JW}}$.) By Theorem 2, there exists a circuit D{D} that implements $\Phi_{\mathrm{JW}}(U_{\tau\sigma^{-1}})$ in depth $O(\log^{2}N)$. Then $C_{2}^{-1}DC_{1}$ is a circuit of depth $O(\log^{2}N)$ transforming from $\phi_{1}$ to $\phi_{2}$. ∎
@@ V Applications
In this section, we discuss applications of our methods to time evolution under widely studied fermionic models, see also Ref. [11] for applications of log-depth routing. We summarize our discussions in Table 2, where we compare to the best previous methods that we are aware of.
Simulating fermionic Hamiltonians is one of the most promising applications of quantum computing, enabling the solution of classically challenging problems in quantum chemistry [30], materials science [31], and high-energy physics [32]. However, in order to simulate fermions with qubit quantum computers, the fermionic operators must be mapped to qubit operators. This leads to a multiplicative overhead in the circuit depth and gate count. In other words, this overhead is defined as the depth or gate count on a conventional qubit quantum computer (see e.g. Refs. [33, 34, 35, 36] for recent experimental demonstrations) divided by the depth or gate count on a fermionic quantum computer that digitally manipulates fermionic particles directly (see e.g. Refs. [37, 38, 39, 40]).
The simplest mapping of fermions to qubits is the Jordan-Wigner encoding [15]. One way to understand this mapping is that it maps products of fermionic operators to products of qubit operators. Naively, this encoding has worst-case depth overhead linear in $N$ for $N$ fermionic sites because some fermionic Hamiltonian terms are mapped to Pauli strings with support on $\Theta(N)$ sites, even if in the fermionic terms only have support on $\Theta(1)$ sites. This is because the Jordan-Wigner encoding labels the fermionic modes along a fictitious one-dimensional chain. For example, the hopping term from one end of the chain to the other has weight $2$ in the fermionic formulation and weight $N$ after mapping to qubits. This is a highly geometrically non-local hopping term, but because of the effective one-dimensionality, even geometrically local hopping terms in low dimension can have high weight if they are not along the direction of the Jordan-Wigner chain. For example, the hopping term in two dimensions which is not along the chain has weight $O(\sqrt{N})$ [34]. While it is widely known that this leads to an at-worst linear depth overhead for fermion time evolution, the reason for this is often miscommunicated: it is not the high weight of the fermion terms which leads to the depth overhead, but the restriction in parallelization. The most straightforward approach for implementing time evolution under a Pauli-string Hamiltonian is done with a so-called Pauli gadget, i.e., the conjugation of a single-qubit rotation with an $N$-qubit CNOT ladder, which requires depth $N$. However, this CNOT ladder can actually be compressed to depth $O(\log N)$ without ancillas [12]. Therefore, a single fermionic Hamiltonian term can be implemented in depth $O(\log N)$ within the Jordan-Wigner encoding, and therefore the depth overhead resulting from the $O(N)$ operator weight is only $O(\log N)$. Instead, the reason why the Jordan-Wigner encoding has $O(N)$ depth overhead for the overall time-evolution task is the parallelization restriction: in the worst case, only one Hamiltonian term can be applied at a time, resulting in a depth scaling with the number of Hamiltonian terms, which is $\Omega(N)$ for a non-trivial Hamiltonian. By contrast, for fermionic quantum computers, there is no such restriction other than the connectivity graph of the Hamiltonian—in most physical Hamiltonians, the connectivity is such that $\Theta(N)$ terms can be implemented in parallel (this is the case in the 2D Fermi-Hubbard model [41]), such that for example $\Theta(N)$ Hamiltonian terms can be implemented with depth $\Theta(1)$.
The Bravyi-Kitaev encoding [14] might at first seem to address this issue [16], as it reduces the operator weight after the fermion-to-qubit mapping to only $O(\log N)$ sites. However, unfortunately, these savings for a single fermionic Hamiltonian term do not carry over to the task of implementing all terms: the Pauli strings to which products of fermion creation and annihilation operators are mapped all have one qubit in common, so they cannot be implemented in parallel [42]. Therefore, in practice, even though the weight reduction can lead to a gate count improvement, for most problems the Bravyi-Kitaev encoding does not show an improvement over the $O(N)$ depth overhead of the Jordan-Wigner encoding. Schemes have been proposed to avoid this restriction by using ancillas [14, 43, 42, 44, 39], which in particular can achieve $O(1)$ overhead for geometrically local models [45, 20]. However, all of these approaches require $\Omega(N)$ ancillas, which might be a prohibitive overhead, especially for fault-tolerant architectures, where qubit count is often more expensive than depth due to the exponential error suppression [46].
Instead, we here rely on the polylog routing schemes from Ref. [11] and this work to reduce the depth of fermion time evolution.
Reducing the fermion-to-qubit overhead in fermion time evolution to permutations
As we will discuss near the end of this section, both of these restrictions are not necessary: this scheme generalizes to any Trotter scheme and any Hamiltonian (without any restrictions on $k$-locality).
On a fermionic quantum computer, which in particular can directly implement $\exp(-itJ_{ij}(a^{\dagger}_{i}a_{j}+\mathrm{h.c.}))$ as a single gate, the minimal depth to implement this scheme depends only on how many of the $J_{ij}$ are nonzero and to what extent the structure of the $J_{ij}$ may be parallelized. For instance, for a full-rank tensor, i.e., with $\Theta(N^{2})$ non-zero terms, the terms can be grouped such that $O(N)$ of them can be implemented in parallel. This means that the minimal depth is $\Omega(N)$. For a non-full-rank tensor, the situation is more complicated: if there are only $\Theta(N)$ non-zero terms, this may still require depth $\Theta(N)$ if all of the terms have one fermion in common. In order to discuss the overhead introduced by the fermion-to-qubit encoding independently of the parallelization restrictions introduced by the Hamiltonian parameters, we define the fermion-to-qubit overhead as the multiplicative overhead in gate depth between an implementation on a fermionic quantum computer and a qubit quantum computer, or in other words, the gate depth on a qubit quantum computer divided by the gate depth on a fermion quantum computer, as we have already defined above. To make this comparison, we assume all-to-all connectivity in both qubit and fermion quantum computers.
In this section, we discuss that, using the schemes introduced in Ref. [11] and this work, fSWAP networks can in fact achieve a worst-case upper bound of $O(\log^{2}N)$ depth overhead compared to fermionic quantum computers for general fermionic Hamiltonians without introducing ancilla qubits. We do so by simply routing the fermion along the Jordan-Wigner chain in-between layers of time evolutions that only act on neighboring fermionic sites, in complete analogy to how routing helps in reducing the depth of all-to-all-connected circuits in 1D nearest-neighbour-connected hardware [8].
Specifically, for our example, to order $t^{2}$,
(We define the product of operators using the following order: $\prod_{i=1}^{N}O_{i}=O_{N}\cdots O_{2}O_{1}$.) This scheme, see also Ref. [11], is a generalization of the time evolution scheme in the original work introducing fSWAP networks [10] and is also closely related to the double factorization scheme [48], where fermion-overhead-free density-density interactions are alternated with fermionic basis transformations. One way to understand the above scheme is that part of the basis transformation in double factorization is absorbed in the Hamiltonian (such that now it is not purely density-density any more), and the rest of the basis transformation is restricted to permutations of fermions.
The terms are grouped into $N_{\mathrm{layers}}$ layers, so that terms within a given layer can be applied in parallel on a fermionic quantum computer. The goal of the permutations is to rearrange the order of the fermion sites in a given layer so that the Hamiltonian terms only act on sites that are adjacent in the Jordan-Wigner encoding, such that $k$-local fermion terms map to $k$-local qubit terms. In the first line of Eq. (14), we insert resolutions of identity in terms of permutation unitaries $U_{\sigma_{l}}$, which apply permutation $\sigma_{l}$ corresponding to layer $l$: $U_{\sigma_{l}}a_{i}U_{\sigma_{l}}^{-1}=a_{\sigma_{l}(i)}$. To enable parallelization of all terms within a given layer in the Jordan-Wigner encoding, the permutation unitaries $U_{\sigma_{l}}$ reorder the fermions along the Jordan-Wigner chain such that, for all $J_{ij}$ within a given layer, the fermions $\sigma_{l}(i)$ and $\sigma_{l}(j)$ are neighboring along the Jordan-Wigner chain. This means that $\exp(-itJ_{ij}(a^{\dagger}_{\sigma_{l}(i)}a_{\sigma_{l}(j)}+\mathrm{h.c.}))\rightarrow\exp(-itJ_{ij}(\Sigma^{+}_{\sigma_{l}(i)}\Sigma^{-}_{\sigma_{l}(j)}+\mathrm{h.c.}))$, where $\Sigma^{\pm}=(X\pm iY)/2$ are qubit Pauli operators, can be implemented in depth $O(1)$ and no parallelization restriction arises from the fermion-to-qubit mapping. Furthermore, for every $l<N_{\mathrm{layers}}$, $U_{\sigma_{l+1}}U_{\sigma_{l}}^{-1}$ can be combined into a single permutation unitary.
As an example, we consider simulation of fermionic Hamiltonians of the form
@@
where $n_{i}=a^{\dagger}_{i}a_{i}$, and $J_{ij},U_{ij}$ are the hopping and interaction coefficient matrices, respectively. This problem is relevant to simulations of materials [31].
@@
The fermionic fast Fourier transform [49] diagonalizes the hopping Hamiltonian provided $J_{ij}$ is translationally invariant on a $d$-dimensional hypercubic lattice with periodic boundary conditions (we focus on the 1-dimensional case here since the $d$-dimensional FFFT is simply the one-dimensional FFFT applied in all dimensions [31], just as for the usual FFT). Concretely,
@@
In the following, we discuss that the FFFT can be applied in polylog depth, implying that a single Trotter step under the hopping Hamiltonian for translationally-invariant systems can be done in polylog depth, as shown in Ref. [11], see also Ref. [49, 39] for the corresponding statement for fermionic quantum computers.
@@
Corollary 3 (Log-depth FFFT with ancillas [11]).
Given any permutation $\sigma\colon[N]\to[N]$, there exists a circuit of depth $O(\log N)$ with $O(N)$ ancillas and measurement and feedforward implementing $\Phi_{\mathrm{JW}}(U_{\sigma})$. In addition, there is a circuit of depth $O(\log N)$ implementing the FFFT using $O(N)$ ancillas and measurement and feedforward.
@@
The materials Hamiltonian—or equivalently, quantum chemistry in the plane-wave dual basis [31]—is essentially a more complicated version of the Fermi-Hubbard model: the coefficients $J_{ij}$ and $U_{ij}$ are full rank, but both are translationally invariant [31]. Therefore, the hopping part can be compressed again to $O(\log^{2}N)$ depth without ancillas and $O(\log N)$ depth with ancillas by using the FFFT and polylog-depth permutations. By contrast, because of the unboundedness of the hoppings, no scheme is known how to achieve this depth in the compact encoding. Instead, the depth in the compact encoding defaults to the same asymptotic depth as using fSWAP networks with transpositions [10] as the same geometric restrictions apply. A scheme to reduce the depth to $O(\log N)$ was proposed in Ref. [39], but required $O(N^{2})$ ancillas. For the interaction term, no scheme is known to apply it in polylog depth without ancillas; however, an $O(\log N)$-depth scheme is known using $O(N\log N)=\tilde{O}(N)$ ancillas [52]. Therefore, the schemes proposed in Ref. [11] and in this work achieve polylog depth for a single Trotter step, as already noted in  [11]. Finally, for obtaining the end-to-end depth for simulating time evolution until time $T$, we need to multiply this by the number of Trotter steps. The best known upper bound is $O(NT)$ from Ref. [53], assumes arbitrarily large Trotter order and fixed total simulation error. Therefore, our schemes achieve $\tilde{O}(NT)$ depth complexity with $\tilde{O}(N)$ qubits. This quasilinear-in-number-of-plane-waves scaling of qubits and depth has to our knowledge not been achieved for qubit quantum computers yet and this scaling is on par (asymptotically) with fermionic quantum computers [39].
(Sub-)routine Compact [17, 20] fSWAP [10] pfSWAP [54] Maskara et al. [11] This work
@@
Fermion permut. NN NN NN 0 11 $N^{2}$ $\log N$ NN $\log^{2}N$ 0
FFFT $\log N NN$ $N\log N$ 0 $\log N$ $N^{2}$ $\log N$ NN $\log^{2}N$ 0
Materials (1 step) NN NN NN 0 $\log N$ $N^{2}$ $\log N$ $N\log N$ $\log^{2}N$ $N\log N$
Materials (full) $^{2} NN$ $N^{2}$ 0 $N\log N$ $N^{2}$ $N\log N$ $N\log N$ $N\log^{2}N$ $N\log N$
2D Fermi-Hubbard 11 NN $\sqrt{N}$ 0 $\log N$ $N^{2}$ $\log N$ NN $\log^{2}N$ 0
Table 2: Asymptotic circuit depth and ancilla count for key fermionic subroutines and applications. All quantities are $O(\cdot)$ upper bounds. Compact refers to any compact encoding, e.g., Verstraete-Cirac [17] or Derby-Klassen [20]; fSWAP is the transposition-based fSWAP network from Ref. [10]; pfSWAP refers to parallel fSWAP from Ref. [39] using $O(N^{2})$ ancillas to brute-force parallelize the CZ circuits in the fSWAP networks; “2D Fermi-Hubbard” applies, more generally, to any model with $O(1)$-range hopping and interactions and assumes periodic boundary conditions for pfSWAP and our work (as we assume the usage of the FFFT for the kinetic term), and the gate counts are for a single Trotter step. The number of Trotter steps $O(N)$ used in “Materials (full)” assumes arbitrarily large Trotter order, fixed evolution time, and fixed simulation error. Maskara et al. [11] makes use of ancillas in the permutation step, whereas our work only uses ancillas for parallelizing the interaction part of the Hamiltonian [52].
@@ VI Discussion and Outlook
In this work, we build on the recent exponential improvement in fermionic routing introduced by Maskara et al. [11], who showed that the worst-case routing overhead in fermion-to-qubit mappings can be reduced from $O(N)$ to $O(\log N)$ by decomposing general fermionic permutations into $O(\log N)$ constant-depth interleaves and using $\Theta(N)$ ancilla qubits and mid-circuit measurements to compress the CZ circuits part of fermi-SWAP networks [10]. We rederive and generalize their results using an alternative construction. Moreover, we show that similar exponential improvements persist even in the ancilla-free setting: any fermionic permutation in the Jordan-Wigner encoding can be implemented in depth $O(\log^{2}N)$ without ancillas. We further extend our analysis to arbitrary product-preserving ternary-tree encodings by constructing efficient mappings between those encodings and the Jordan-Wigner encoding in $O(\log^{2}N)$ depth. This establishes that fermionic routing can be efficiently performed in a broad class of fermionic encodings. As a result, key primitives such as the fermionic fast Fourier transform (FFFT) and single-Trotter-step material simulations can be executed in polylogarithmic depth using only $\tilde{O}(N)$ qubits for $N$ fermionic modes.
Our results assume hardware with all-to-all connectivity, but they also imply upper bounds for limited-connectivity devices. In particular, the worst-case overhead for fermionic simulation in constrained architectures is $O(\log^{2}N)$ times the worst-case depth required to route qubits on the underlying hardware graph, if that hardware graph has $\Omega(N)$ disjoint edges [55]. For instance, qubit systems with grid connectivity can implement the FFFT in depth $O(\sqrt{N}\log^{2}N)$; current neutral-atom architectures achieve $O(\sqrt{N}\log^{3}N)$ depth, while upgraded neutral-atom hardware could reach $O(\log^{3}N)$ depth [8]. This represents an exponential improvement over the prior best-known FFFT algorithm, which achieved $O(N\log N)$ depth [31] on a linear geometry, with no mechanism to exploit higher connectivity. The above-cited depths in the presence of hardware constraints is likely too pessimistic as it neglects additional structure in the circuits. In fact, the FFFT is an example where that structure can be exploited, and Maskara et al. [11] showed several others.
Finally, these results motivate further study of fermionic routing and simulation on hardware with restricted connectivity, particularly the trade-offs among ancilla count, mid-circuit measurement and feedforward, and depth overhead. Beyond asymptotic scaling, a detailed analysis of constant factors in depth and gate count will be essential for identifying the most practical regimes of advantage. It also remains an open question whether analogous circuit-structural insights could lead to improved boson-to-qubit encodings [56, 57, 58].